OpenStudy (anonymous):

hey how do i find dy/dx of tan(X+y)=4x?

OpenStudy (anonymous):

calc 1 ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

alright and have you learned implicit differentiation

OpenStudy (anonymous):

yea im ok with it

OpenStudy (anonymous):

i wud subtrac 4x first right?

OpenStudy (anonymous):

um not necessarily you can just simply take the dy/dx of both sides right now if you wanted to

OpenStudy (anonymous):

oh ,well ill subtract since our teach showed us that n hell b grading it lol

OpenStudy (anonymous):

alright now differentiate with the respects to x

OpenStudy (anonymous):

dy/dx is respect to x or y?

OpenStudy (anonymous):

nvm that, what do i do with the tan(x+y) ?

OpenStudy (anonymous):

is it sec^2(x+Y)?

OpenStudy (anonymous):

nvm that, what do i do with the tan(x+y) ?

OpenStudy (anonymous):

use the chain rule which would be sec^2(u)(u')

OpenStudy (anonymous):

deriv of y =1 right?

OpenStudy (anonymous):

k so got sec^2(x+y)=4 now

OpenStudy (anonymous):

\[\frac{d}{dx}[y]=\frac{dy}{dx}\]

OpenStudy (anonymous):

in respects to x... you can't find the derivative of y since it's not x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so you should get \[\sec^2(x+y)(1+\frac{dy}{dx})=4\]

OpenStudy (anonymous):

ahh ok

OpenStudy (anonymous):

alright so what do you get for an answer

OpenStudy (anonymous):

one sec so now i factor out dy /dx but what happens to sec?

OpenStudy (anonymous):

oh wait is answer 4/sec^2(X+y)?

OpenStudy (anonymous):

multiply the two to get \[\sec^2(x+y)+\frac{dy}{dx}\sec^2(x+y)\] - sec^2(x+y and divide by sec^2(x+y) \[\frac{dy}{dx}=\frac{4-\sec^2(x+y)}{\sec^2(x+y)}\]

OpenStudy (anonymous):

thanks so much